A matchbox measures 4cm multiple 2.5 cm multiple 1.5 cm multiple. What will be the volume of the packet cointaing 12 such boxes?
 

Grade:9

1 Answers

Vikas Amritiya
2 years ago
Since the matchbox is cuboidal in shape, the volume of each matchbox will be equal to the volume of thecuboid.

Thevolume of the cuboidof lengthl, breadthb, and heighth, V = l × b × h

Length of the matchbox, l = 4 cm, Breadth of the matchbox, b = 2.5 cm, Height of the matchbox, h = 1.5 cm

The volume of each matchbox = l × b × h

= 4 cm × 2.5 cm × 1.5 cm

= 15 cm3

Volume of 12 matchboxes = 12 × 15cm3= 180cm3

Thus, the volume of thepacket containing 12 such matchboxes is 180 cm3.